Evaluate C∫ (2x2 + 3y2)dx, where C is the curve given by x(t)=at2 , y(t)=2at, 0≤t≤1.


1
Expert's answer
2021-06-09T15:10:30-0400

Let us evaluate "\int_C (2x^2 + 3y^2)dx", where "C" is the curve given by "x(t)=at^2 , y(t)=2at, 0\u2264t\u22641."


"\int_C (2x^2 + 3y^2)dx=\int_0^1 (2(at^2)^2 + 3(2at)^2)d(at^2)=\int_0^1 (2a^2t^4 + 12a^2t^2)2atdt=
\int_0^1 (4a^3t^5 + 24a^3t^3)dt=4a^3\int_0^1 (t^5 + 6t^3)dt=4a^3(\frac{t^6}{6}+6\frac{t^4}{4})|_0^1=
4a^3(\frac{1}{6}+\frac{3}{2})=4a^3\frac{10}{6}=\frac{20a^3}{3}."